OpenStudy (anonymous):

sin^2(x)+cos(x)+1=0, 0<=x<2pi

OpenStudy (goformit100):

Plz Post the full Question

OpenStudy (anonymous):

start with \((1-\cos^2(x))+\cos(x)+1=0\) and solve the resulting quadratic equation

OpenStudy (anonymous):

I get x = pi and x = undefined, is this right?